# Adversarial Style Augmentation via Large Language Modelfor Robust Fake News Detection

### v2

@@ Abstract
The spread of fake news negatively impacts individuals and is regarded as a significant social challenge that needs to be addressed. A number of algorithmic and insightful features have been identified for detecting fake news. However, with the recent LLMs and their advanced generation capabilities, many of the detectable features (e.g., style-conversion attacks) can be altered, making it more challenging to distinguish from real news. This study proposes adversarial style augmentation, AdStyle, to train a fake news detector that remains robust against various style-conversion attacks. Our model’s key mechanism is the careful use of LLMs to automatically generate a diverse yet coherent range of style-conversion attack prompts. This improves the generation of prompts that are particularly difficult for the detector to handle. Experiments show that our augmentation strategy improves robustness and detection performance when tested on fake news benchmark datasets.
Adversarial Style Augmentation via Large Language Modelfor Robust Fake News Detection
Sungwon Park1, Sungwon Han1, Meeyoung Cha2 1School of Computing, KAIST, Daejeon, Republic of Korea 2MPI for Security and Privacy, Bochum, Germany {psw0416, lion4151}@kaist.ac.kr, mia.cha@mpi-sp.org
1 Introduction
With the widespread use of the internet and the emergence of various social media platforms, people have begun to freely share information they know and their own creative stories. However, the ease of sharing and consuming information beyond traditional news organizations has also significantly contributed to the spread of fake news Cha et al. (2021); Fisher et al. (2016); Treen et al. (2020). Fake news is created to provide readers with false information as propaganda or to guide public perception in a desired direction for intentional objectives Lazer et al. (2018); Yi et al. (2024). The spread of such false information has had profound negative impacts on individuals and society, becoming a major social challenge that needs to be addressed.
Manually determining the authenticity of all information on the internet is extremely costly in terms of time and resources. Thus, prior advances have mainly focused on developing automated fake news detectors using machine learning techniques Pérez-Rosas et al. (2018); Reis et al. (2019); Shu et al. (2019). For example, several studies have extracted sentiment-related or political features commonly found in fake news, either through human-crafted Potthast et al. (2018); Rashkin et al. (2017) or data-driven methods Reis et al. (2019); Shu et al. (2019), and used them for detection. With the advent of language models (e.g., GPT, BERT) Kenton and Toutanova (2019); Radford et al. (2019), efforts have been made to learn the textual style of fake news by extracting sentence embeddings to train detectors Chen et al. (2021); Yang et al. (2022).
While these extracted textual styles or human-crafted features have proven effective in identifying fake news, they allow attackers to bypass detection. In particular, the advent of large language models (LLMs) Chowdhery et al. (2022); OpenAI (2023) has made it possible to easily and automatically paraphrase sentences in a user-desired direction through prompts (e.g., "change the given text to an objective and professional style"), making it more difficult to distinguish between AI-generated fake news and real news (i.e., style-conversion attacks) Koenders et al. (2021); Zhou et al. (2019). Recent literature has proposed strategies to augment the styles of input text via some manually defined style-conversion prompts, which are taken by LLM for paraphrasing, but it remains challenging to address all types of prompts that attackers might use to deceive detectors Wu and Hooi (2023).
@@ 1 Introduction
In this study, we propose adversarial style augmentation, AdStyle, to train a robust fake news detector that can withstand various style-conversion attacks by attackers. In contrast to earlier work, which used predefined style-conversion prompts as detector-agnostic augmentations, AdStyle tries to find prompts for adversarial style augmentations that are specific to the detector, which makes predictions uncertain. This means that our augmentations add noise to the style features in the direction of the detector’s decision boundary, while maintaining the text’s content integrity. Specifically, to search for and optimize prompts for LLM that are adversarial to the detector, we introduce an automated prompt engineering technique Yang et al. (2023). By providing LLMs with style-conversion prompts and the detector’s performance under these augmentations, LLM can infer patterns between the prompts and performance, enabling the search for the most adversarial prompts.
AdStyle proceeds as follows: To generate augmented samples, we first select a random subset of the dataset and apply a pool of style-conversion prompts to it. Then, each augmentation set created by different prompts is fed into the detector, and the score based on the AUC between the predictions and the ground-truth labels is measured. The prompt-score pairs are provided to the LLM, which uses this information to generate new style-conversion prompt candidates. From these candidates, we select the top-$k$ prompts that are diverse and make the detector’s predictions most uncertain without significantly altering the original content. These selected prompts are applied to the entire dataset and used for detector training. The chosen prompts are added to the prompt pool, which is repeated over several training rounds.
We experimented with existing fake news benchmark datasets under various style-conversion attack scenarios. As a result, our augmentation strategy demonstrated higher robustness and detection performance compared to previous methodologies. Additionally, our approach is scalable to different attack strategies by adding new attack prompts to the style-conversion prompt pool. We plan to release our code after publication.
2 Related Work
2.1 Automated Detection of Fake News
@@
Manually detecting fake news among the vast amount of content on the internet is intractable. Consequently, many studies have focused on leveraging machine learning techniques to automatically detect fake news Pérez-Rosas et al. (2018); Reis et al. (2019); Shu et al. (2019). For example, supervised learning methodologies extract textual features from fake news texts using benchmark datasets and ground-truth labels Reis et al. (2019). These textual features can include deep features based on artificial neural networks Reis et al. (2019); Shu et al. (2019) as well as manually defined features such as sentiment and political bias Potthast et al. (2018); Rashkin et al. (2017). Other approaches include domain adaptation methods to enhance detection generalizability across various domains and topics Mosallanezhad et al. (2022); Nan et al. (2022), and knowledge-based methods that rely on external data to distinguish false information Dun et al. (2021). With the advent of LLMs, new methods have emerged that utilize LLMs’ prior knowledge to identify fake news Hu et al. (2024), including research on detecting machine-generated fake news Chen and Shu (2023).
In this work, we aim to prevent malicious attackers from deceiving automated fake news detectors through text modifications such as paraphrasing. The proposed method is an augmentation strategy to enhance robustness against text style perturbations as an add-on to existing detection models. Our contribution is agnostic to the design of the detection model.
Figure 1: Illustration of how AdStyle works. Our model consists of multiple rounds of training. In each round, the model uses the style-conversion prompt and prediction confusion score from the previous round to generate prompt candidates that can maximize the confusion of the detector model. Subsequently, considering aspects such as adversarialness, coherency, and diversity, a subset of the training dataset is used to select the top-$k$ prompts, which are finally used as augmentations in training.
2.2 Attack on Fake News Detection
@@
To test the robustness of fake news detection, various attack methods have been studied Koenders et al. (2021); Zhou et al. (2019). These include injecting misinformation by changing the order of subjects and objects or causes and effects while maintaining the textual features used for detection Koenders et al. (2021). Other methods involve creating fact distortions by altering or exaggerating words related to people, time, or places while preserving the sentence structure Zhou et al. (2019). Additionally, approaches that use the text generation capabilities of LLMs to change the style of sentences have been proposed Wu and Hooi (2023).
Our method aims to perform adversarial training of the detection model by generating augmentations that perturb the text’s style while preserving the content as much as possible. This approach reduces the impact of spurious style features on the detection model, making it more robust against such attacks.
2.3 Prompt Engineering for LLM
Training with extensive text corpora, LLMs have demonstrated their utility across various domain tasks Chowdhery et al. (2022); OpenAI (2023); Pang et al. (2024). To better harness the prior knowledge and reasoning abilities of LLMs, strategies for providing appropriate input prompts have also been researched. For instance, in-context learning methods involve providing examples of the desired input-output format to guide the LLM in producing the correct output Dong et al. (2022). Additionally, engineering techniques such as chain-of-thought prompting, which includes additional reasoning steps in the responses, have been developed to enhance the reasoning abilities of LLMs Wei et al. (2022). Recently, methods for automatically finding the most suitable prompts for a given task, known as automated prompt engineering, have shown promising results Yang et al. (2023); Zhou et al. (2022).
In this paper, we adopt automated prompt engineering techniques to identify adversarial prompts that most effectively confuse the fake news detector. This process replaces the traditional gradient descent method used in the image domain to find adversarial noise, enabling the creation of adversarially augmented versions of input text via LLM.
3 Method
3.1 Overview
@@
Let $\mathcal{D}=\{(\mathbf{d}_{i},y_{i})\}_{i=1}^{N}$ be a dataset containing news $\mathbf{d}_{i}$ and the corresponding ground-truth binary veracity label $y_{i}$ (indicating whether the news is true or fake). Each news item $\mathbf{d}_{i}$ is composed of natural language-based text. This study aims to train a language model based fake news detector $f$ using the labeled dataset to predict the veracity labels. Our main goal is to develop a detector that remains robust even when an attacker perturbs the textual style, such as the order and format, while preserving the meaning of the sentences.
@@
Figure 1 illustrates how our model works. AdStyle generates style-conversion prompts and performs augmentation over multiple rounds. Each round includes a following process. Firstly, leveraging the reasoning ability of LLM, adversarial style-conversion prompt candidates are generated. These style-conversion prompts contain instructions on how to transform the given sentences and are used as inputs to the LLM along with the original sentences for conversion. Inspired by automated prompt engineering Yang et al. (2023), the style-conversion prompts and detector prediction score pairs used in previous rounds are included as in-context demonstrations to guide the LLM in searching prompt candidates that maximally confuse the detector (Section 3.2). Next, a subset of the dataset is selected, and these discovered candidates are applied to perform conversions. The converted samples are then evaluated to determine how much they confuse the detector. From these candidates, the top-$k$ prompts that are diverse and maintain the original content’s meaning while most effectively confusing the detector are selected (Section 3.3). These selected prompts are used as an augmentation method to train the detection model in the current round. We describe each step’s details below.
3.2 Generating Adversarial Style-Conversion Prompts with LLM
The style-conversion prompts we aim to generate are instructions that perturb only the textual style, such as the structure or format of the sentence, while preserving the content of the given input text. For example, an instruction like “Rewrite the following article in an objective and professional tone” can be used to transform the stylistic features of fake news to resemble real news. However, if such instructions are heuristically defined, they may not align with the actual decision boundary of the detector, reducing the training efficiency. Furthermore, if the instructions remain fixed throughout the training, the detector may memorize these conversion patterns, resulting in overfitting.
Therefore, instead of predefining and fixing the instructions for conversion, we aim to find adversarial conversion prompts that add noise in the direction of the decision boundary of current state of detector, maximizing the confusion in the detector’s predictions. This approach is similar to adversarial training commonly used in the computer vision domain Madry et al. (2018), which enhances robustness against slight perturbations in the input. However, for textual data, the discrete nature of the input makes it challenging to add noise directly through the detector’s gradient as in the image domain. To address this, we measure the confusion in prediction by conversion prompts (i.e., prediction confusion score), and feed this prompt-score pair to LLM for generating adversarial prompts, motivated from automated prompt engineering techniques.
@@ 3.2 Generating Adversarial Style-Conversion Prompts with LLM
Now you will help me minimize a fake news detector performance value with style transfer attack instruction. I have some style transfer attack instructions and the fake news detector performance value using those instructions. The pairs are arranged in descending order based on their function values, where lower values are better. Input: [Style-Conversion Prompt Example] Score: [Prediction Confusion Score] Input: … … Give me a new style transfer attack instruction that is different from all pairs above, and has a performance value lower than any of the above.
@@
Figure 2: Prompt for generating adversarial style-conversion prompts. The text in blue represents the score trajectory component, while the remaining text represents the problem description component.
@@
Figure 2 shows an example of the LLM input used to generate adversarial style-conversion prompts. The LLM input consists of the problem description and the score trajectory components.
Problem description component.
@@
This component includes the problem description, the objective, and constraints on the response necessary for generating style-conversion prompts. For example, a sentence like “Minimize a fake news detector performance value with style transfer attack instruction” informs the LLM of the intent behind the conversion prompt.
Score trajectory component.
@@
Previous work has shown that LLMs can learn patterns from in-context demonstrations provided as input Dong et al. (2022); Yang et al. (2023). This component leverages this ability by providing previous round’s style-conversion prompts and their corresponding prediction confusion scores in the form of in-context demonstrations [^1] In the first round, a predefined set of prompts is used.. The score is measured by selecting a subset $\mathcal{B}$ from the entire training dataset $\mathcal{D}=\{(\mathbf{d}_{i},y_{i})\}_{i=1}^{N}$, applying a conversion prompt $c$ to create a new set $\mathcal{B}^{c}=\{(\mathbf{d}^{c}_{i},y_{i})\}_{i=1}^{M}$, where $\mathbf{d}^{c}_{i}=\text{Convert}(c,\mathbf{d}_{i}),N>>M$, and then measuring the AUC score between the predictions and the ground-truth labels when $\mathcal{B}^{c}$ is fed into the detector. Specifically, the score of a conversion prompt $s_{c}$ is defined as:
$s_{c}=|0.5-\text{AUC}(\{y_{i}\}_{i=1}^{M},\{f(\mathbf{d}^{c}_{i})\}_{i=1}^{M})|.$ (1)
@@
A lower score indicates a higher level of prediction confusion, implying that the conversion prompt has caused the detector’s predictions to become random with respect to the labels. Finding conversion prompts that cause high confusion (i.e., low score) suggests that the detector has not yet learned to handle those stylistic features, and the conversion prompts have placed the samples near the detector’s decision boundary, making them difficult to distinguish. By showing the LLM these in-context demonstrations, it can generate conversion prompts that differ from past ones while maximizing confusion (i.e., minimizing the score). Note that we avoid selecting conversion prompts that flip the detector’s original predictions (i.e., AUC <<much-less-than$<$ < 0.5), as these often disrupt the content along with the stylistic features, which is undesirable. We extract $S$ style-conversion prompts at a time using the input described above.
3.3 Selecting Top-$k$ Adversarial Prompts

### v3

@@ Abstract.
The spread of fake news harms individuals and presents a critical social challenge that must be addressed. Although numerous algorithmic and insightful features have been developed to detect fake news, many of these features can be manipulated with style-conversion attacks, especially with the emergence of advanced language models, making it more difficult to differentiate from genuine news. This study proposes adversarial style augmentation, AdStyle, designed to train a fake news detector that remains robust against various style-conversion attacks. The primary mechanism involves the strategic use of LLMs to automatically generate a diverse and coherent array of style-conversion attack prompts, enhancing the generation of particularly challenging prompts for the detector. Experiments indicate that our augmentation strategy significantly improves robustness and detection performance when evaluated on fake news benchmark datasets.
Misinformation, Adversarial Training, Large Language Model
copyright: acmlicensedjournalyear: 2025conference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714569isbn: 979-8-4007-1274-6/25/04ccs: Security and privacy Software and application securityccs: Computing methodologies Artificial intelligence
1. Introduction
In today’s digital landscape, people seek information through various channels, including social media. News content continues to hold substantial importance on these platforms, significantly shaping individuals’ thoughts and decisions. However, the ease of sharing and consuming information outside traditional news outlets has introduced several challenges. Among them is the rise of alternative media and news-like content, which can contain completely or partially false information (Vosoughi et al., 2018). The spread of such misinformation has negative impacts on individuals and is considered a major social challenge that requires urgent attention (Cha et al., 2021).
Given the vast amount of information shared on social platforms, manually detecting fake news has become impractical. Consequently, these efforts are increasingly supported and managed by algorithms. For example, sentiment and topic features, as well as the temporal and structural patterns of diffusion networks, have proven effective in detecting fake news (Kwon et al., 2013; Potthast et al., 2018; Rashkin et al., 2017). These patterns can be utilized by machine learning and deep learning (Ma et al., 2016; Shu et al., 2019; Pérez-Rosas et al., 2018). Large language models (LLMs) have also been used to learn the textual style of fake news by extracting sentence embeddings to train detectors (Augenstein et al., 2024b; Hu et al., 2024; Chen and Shu, 2023).
Although certain textual styles or linguistic cues can indicate fake news, attackers can circumvent such detection by rearranging the order of subjects and objects or by imitating specific styles (Koenders et al., 2021; Zhou et al., 2019). LLMs can easily paraphrase sentences in any desired manner through prompts (e.g., ”change this text into a NYTimes style”), making it challenging to differentiate AI-generated news from authentic ones. These manipulations, referred to as style-conversion attacks, significantly complicate the task of maintaining information veracity in the digital media landscape (Koenders et al., 2021; Zhou et al., 2019).
@@ 1. Introduction
In this study, we present AdStyle, an adversarial style augmentation method designed to train a fake news detector that can withstand various style-conversion attacks. Unlike previous methods that relied on predefined style-conversion prompts, our approach identifies prompts to generate adversarial style augmentations tailored to a specific detector. This process adds noise to the style features within the detector’s decision boundary, while preserving content-wise integrity. To customize prompts that are adversarial to the detector, we implement an automated prompt engineering technique, similar to (Yang et al., 2023). By providing style-conversion prompts and evaluating the detector’s performance under these augmentations, the LLM can discern patterns between prompts and performance, facilitating search for the most effective prompts.
When evaluated on fake news benchmarks such as PolitiFact, GossipCop, and Constraint under various style-conversion attack scenarios, our augmentation strategy exhibited enhanced robustness and detection performance compared to current state-of-the-art methods. Our strategy preserves the content of the sentence while altering its structure to increase perplexity according to the LLM-based detector. This adjustment results in a sentence structure that the LLM has not frequently encountered during its pre-training phase, effectively deterring malicious style-conversion attacks, such as paraphrasing. Additionally, our method serves as an augmentation strategy that can be integrated with any existing detection models, regardless of their design.
We release the code and implementation details of our fake news detection against style-conversion attacks for the broader use and greater impact within the research community and industry: https://github.com/deu30303/AdStyle.
2. Related Work
2.1. Automated Detection of Fake News
@@
Among the machine-driven fake news detection methods are learning methodologies extract textual features from fake news texts using benchmark datasets and ground-truth labels (Reis et al., 2019). These textual features can include deep features based on artificial neural networks (Shu et al., 2019) as well as manually defined features such as sentiment and political bias (Kwon et al., 2013; Potthast et al., 2018; Rashkin et al., 2017). Other approaches include domain adaptation methods to enhance detection generalizability across various domains and topics (Mosallanezhad et al., 2022; Nan et al., 2022), and knowledge-based methods that rely on external data to distinguish false information (Dun et al., 2021). With the advent of LLMs, new methods have emerged that utilize LLMs’ prior knowledge to identify fake news (Hu et al., 2024), including research on detecting machine-generated fake news (Chen and Shu, 2023). One approach involves using LLMs to generate synthetic reactions and comments on news articles from diverse perspectives, which are then leveraged for fake news detection (Wan et al., 2024). Furthermore, LLM-generated logic has been employed to enhance the interpretability of fake news detection models (Liu et al., 2024). Despite these advancements, fake news detection using LLMs remains a challenging task due to factuality hallucinations, which can undermine reliability (Augenstein et al., 2024a).
2.2. Attack on Fake News Detection
@@
Various attack methods have been studied to test the robustness of fake news detection (Koenders et al., 2021; Zhou et al., 2019). These methods include injecting misinformation by rearranging subjects and objects or causes and effects while retaining the textual features used for detection (Koenders et al., 2021). Other techniques involve creating fact distortions by modifying or exaggerating words related to people, time, or places, while preserving sentence structure (Zhou et al., 2019). Another line of research simulates the behavior of malicious users on social media by manipulating news article comments or employing a multi-agent reinforcement learning framework to generate adversarial content (Le et al., 2020; Wang et al., 2023). Recently, methods leveraging the text generation capabilities of LLMs to modify the writing style of sentences have also been proposed (Wu and Hooi, 2023). In this study, we propose an adversarial training method to mitigate style-conversion attacks in fake news detection. Our approach utilizes automated prompt engineering to identify the most effective style-conversion prompts for a given fake news dataset (Yang et al., 2023; Zhou et al., 2022).
Figure 1. Illustration of AdStyle. The model undergoes multiple rounds of training. In each round, it utilizes the style-conversion prompt and prediction confusion score from the previous round to create prompt candidates aimed at maximizing the detector model’s confusion. Subsequently, a subset of the training dataset is used to select the top-$k$ prompts, based on factors such as adversarialness, coherency, and diversity. These selected prompts are then employed as augmentations in the training process. Repetition of this process allows the model to consider a wide range of adversarial prompts, making the director more resilient to future attacks.
3. Method
3.1. Overview
@@
Let $\mathcal{D}=\{(\mathbf{d}_{i},y_{i})\}_{i=1}^{N}$ be a dataset containing news $\mathbf{d}_{i}$ and the corresponding ground-truth binary veracity label $y_{i}$ (indicating whether the news is true or fake). Each news item $\mathbf{d}_{i}$ is composed of natural language-based text. Our goal is to train a fake news detector $f$ based on the labeled dataset and predict veracity labels. In doing so, we ensure that a detector remains robust even when an attacker perturbs the textual style, such as the order and format, while preserving the meaning of the sentences.
@@
Figure 1 illustrates the workflow. Our method, called AdStyle, generates style-conversion prompts and performs augmentation over multiple rounds of the following process. Firstly, leveraging the reasoning ability of LLM, adversarial style-conversion prompt candidates are generated. These style-conversion prompts contain instructions on how to transform the given sentences and are used as inputs to the LLM along with the original sentences for conversion. Following the automated prompt engineering (Yang et al., 2023), the style-conversion prompts and detector prediction score pairs used in previous rounds are included as in-context demonstrations to guide the LLM in searching prompt candidates that maximally confuse the detector (Section 3.2). Next, a subset of the dataset is selected, and these discovered candidates are applied to perform conversions. The converted samples are then evaluated to determine how much they confuse the detector. From these candidates, the top-$k$ prompts that are diverse and maintain the original content’s meaning while most effectively confusing the detector are selected (Section 3.3). These selected prompts are used as an augmentation method to train the detection model in the current round. We describe each step’s details below.
@@
Now you will help me minimize a fake news detector performance value with style transfer attack instructions. I have a set of style transfer attack instructions that I will use to minimize the fake news detector’s performance. The pairs are arranged in descending order based on their function values, where lower values are better.
Input: [Style-Conversion Prompt Example]
Score: [Prediction Confusion Score]
Input: … …
Give me a new style transfer attack instruction that is different from all pairs above and has a lower performance value.
@@
Figure 2. Prompt for generating adversarial style-conversion prompts. The text in blue represents the score trajectory, and the remaining text represents the problem description.
3.2. Adversarial Style-Conversion Prompts
The style-conversion prompts we generate are instructions that perturb only the textual style, such as the structure or format of the sentence, while preserving the content of the given input text. For example, an instruction like “Rewrite the following article in an objective and professional tone” can transform the stylistic features of fake news to resemble authentic news. However, if these instructions are heuristically defined, they may not align with the detector’s actual decision boundary, thereby reducing training efficiency. Furthermore, if the instructions remain fixed throughout the training process, the detector may memorize these conversion patterns, leading to overfitting.
Instead of predefining and fixing the instructions for conversion, we identify adversarial conversion prompts that introduce noise directed toward the decision boundary of the current detector, thereby maximizing prediction confusion. This approach is similar to adversarial training commonly used in the computer vision domain (Madry et al., 2018), which enhances robustness against slight input perturbations. However, for textual data, the discrete nature of the input makes it challenging to add noise directly through the detector’s gradient as in the image domain. To address this, we measure prediction confusion using conversion prompts (i.e., prediction confusion score) and feed this prompt-score pair to the LLM for generating adversarial prompts, inspired by automated prompt engineering techniques.
@@ 3.2. Adversarial Style-Conversion Prompts
Figure 2 shows an example of the LLM input used to generate adversarial style-conversion prompts. The LLM input consists of the problem description and the score trajectory components. We describe each component in turn.
@@
Problem description component. This component includes the problem description, the objective, and constraints on the response necessary for generating style-conversion prompts. For example, a sentence like “Minimize a fake news detector performance value with style transfer attack instruction” informs the LLM of the intent behind the conversion prompt.
@@
Score trajectory component. Previous research has demonstrated that LLMs can learn patterns from in-context demonstrations provided as input (Dong et al., 2022; Yang et al., 2023). The score trajectory component leverages this ability by providing style-conversion prompts from previous rounds and their corresponding prediction confusion scores in the form of in-context demonstrations. In the first round, a predefined set of prompts is excerpted from (Wu and Hooi, 2023). The score is calculated by selecting a subset $\mathcal{B}$ from the entire training dataset $\mathcal{D}=\{(\mathbf{d}_{i},y_{i})\}_{i=1}^{N}$, applying a conversion prompt $c$ to create a new set $\mathcal{B}^{c}=\{(\mathbf{d}^{c}_{i},y_{i})\}_{i=1}^{M}$, where $\mathbf{d}^{c}_{i}=\text{Convert}(c,\mathbf{d}_{i}),N>>M$, and then measuring the AUC score between the predictions and the ground-truth labels when $\mathcal{B}^{c}$ is fed into the detector. Specifically, the score of a conversion prompt $s_{c}$ is defined as:
(Eq. 1) $s_{c}=|0.5-\text{AUC}(\{y_{i}\}_{i=1}^{M},\{f(\mathbf{d}^{c}_{i})\}_{i=1}^{M})|.$
@@
A lower score indicates a higher level of prediction confusion, implying that the conversion prompt has caused the detector’s predictions to become random with respect to the labels. Finding conversion prompts that lead to high confusion (i.e., low score) suggests that the detector has not yet learned to handle those stylistic features, and the conversion prompts have placed the samples near the detector’s decision boundary, making them difficult to distinguish. By providing these in-context demonstrations, the LLM can generate conversion prompts that differ from previous ones while maximizing confusion (i.e., minimizing the score). It is important to avoid selecting conversion prompts that flip the detector’s original predictions (i.e., AUC <<much-less-than$<$ < 0.5), as such cases will disrupt the content and stylistic features. We extract $S$ style-conversion prompts at a time using the input described above.
3.3. Selecting Top-$k$ Adversarial Prompts